OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). A. y = 2x + 6 B. y = −1/2x + 3 C. y = −1/2x + 6 D. y = 2x + 3

OpenStudy (anonymous):

La pendiente de la recta perpendicular es -1/2, porque (2)(-1/2)=-1 \[−\frac{1}{2}=\frac{y−3}{x−6} \\y=6−\frac{x}{2}\] La respuesta es la C.

OpenStudy (anonymous):

gracias!